# Shoal++: High Throughput DAG BFT Can Be Fast!

### v1

@@ Abstract.
Today’s practical partially synchronous Byzantine Fault Tolerant (BFT) consensus protocols trade off low latency and high throughput. On the one end, traditional BFT protocols such as PBFT and its derivatives optimize for latency. They require, in fault-free executions, only 3 message exchanges to commit, the optimum for BFT consensus. However, this class of protocols typically relies on a single leader, hampering throughput scalability. On the other end, a new class of so-called DAG-BFT protocols demonstrates how to achieve highly scalable throughput by separating data dissemination from consensus, and using every replica as proposer. Unfortunately, existing DAG-BFT protocols pay a steep latency premium, requiring on average 10.5 message exchanges to commit a transactions.
@@
This work aims to soften this tension, and proposes Shoal++, a novel DAG-based BFT consensus system that offers the throughput of DAGs while reducing commit latency to an average of 4.5 message exchanges. Our empirical findings are encouraging, showing that Shoal++ achieves throughput comparable to state-of-the-art DAG BFT solutions while reducing latency by up to 60%.
submissionid: 688copyright: none
1. Introduction
@@
This paper presents Shoal++, a novel partially synchronous DAG-based Byzantine Fault Tolerant (BFT) consensus protocol that matches the throughput of contemporary DAG-BFT protocols while significantly reducing latency, closing the gap with the theoretical optimum.
@@
BFT consensus offers the attractive abstraction of a single, always available server, that remains correct even as a subset of replicas may fail arbitrarily. Fueled by recent advances to simplicity and practicality (Castro and Liskov, 2002; Yin et al., 2019; Danezis et al., 2022; Spiegelman et al., 2022a), interest and adoption of BFT protocols is skyrocketing: it is at the core of aspiring multi-national projects such as the digital Euro (dig, [n. d.]), confidential data sharing frameworks (ccf, [n. d.]) and the growing Web3 industry. Already today, numerous blockchain companies (Team, 2022, 2019, 2020, 2023) are actively deploying BFT protocols (Spiegelman et al., 2022a, 2023; Chen and Micali, 2019; Rocket, [n. d.]; Gelashvili et al., 2022), supporting millions of daily worldwide users.
@@
Following the seminal PBFT (Castro and Liskov, 2002) protocol, most traditional BFT consensus protocols (Yin et al., 2019; Kotla et al., 2007; Gueta et al., 2019) operate in a primary-backup like regime in which a single, dedicated leader replica proposes an ordering of commands, and follower replicas vote to ensure consistency and durability.
@@
As demonstrated by PBFT, this approach can enjoy excellent latency in fault-free executions, requiring only three message exchanges (henceforth message delays, or md) to order a leader’s proposal, the optimum achievable for BFT protocols. A dedicated leader, however, poses not only a single point of failure, but constrains the achievable throughput to the processing and networking bandwidth available to a single server. To improve scalability, and meet the high throughput demands of modern applications, recent proposals emphasize maximizing resource utilization and balancing load across all replicas.
[…]
@@
State-of-the-art DAG protocols (Spiegelman et al., 2022a, 2023) avoid running an explicit consensus protocol, and instead embed the consensus process onto the DAG directly: they ensure that all replicas converge on the same view of the DAG, and interpret it’s structure locally to implicitly determine commitment. Consensus simulates leaders by assigning pre-determined ”fixpoint” nodes in the DAG (henceforth anchors) and confirms agreement of an anchor by waiting until it has gathered sufficient references in future DAG rounds that guarantee its durability.
@@
Upon committing an anchor node, all of its causally referenced proposals are implicitly committed and ordered. Such design is appealingly simple – it requires no additional communication for consensus and avoids the notoriously complex view change logic inherent to traditional BFT protocols. It is also very resource efficient – each replica acts as a proposer, thus allowing high throughput with a large number of replicas. As a result, DAG-BFT has prompted swift adoption by several large blockchain systems (Team, 2023, 2022).
@@
However, despite their desirable features, today’s DAG protocols fall short in terms of latency. They require several rounds, each consisting of a RB, in order to commit a transaction. Bullshark (Spiegelman et al., 2022a) for instance requires, even in fault-free cases, up to 4 rounds to commit a proposal (12md in expectation, §3.2) ; recent advances made by Shoal (Spiegelman et al., 2023) reduce it to 3 rounds (10.5md expected, §3.2). A far cry from the 3md possible by traditional BFT protocols!
This work argues that these latency shortcomings is not fundamental to DAG based BFT, and proposes a new DAG-based BFT system Shoal++ that reduces the latency gap close to the optimum.
@@
We begin by analyzing sources of end-to-end (e2e) transaction latency in DAG protocols (§3.2); at a high level, it can be broken down into three stages. (i) DAG proposals happen only at regular intervals (rounds), and transactions that narrowly miss inclusion (e.g. arrive just after a batch is formed) must wait for the next round. We refer to the time a transaction must wait to be included in a DAG proposal as Queuing Latency. (ii) In order to be committed, a proposed DAG node must be referenced by a committed anchor. Unfortunately, in existing DAG protocols (Spiegelman et al., 2022a, 2023), at most one anchor is scheduled per round [^1] Bullshark (Spiegelman et al., 2022a) schedules an anchor every second round; Shoal (Spiegelman et al., 2023) improves it to every round. This results in additional Anchoring latency, the time until a node is referenced by the next committed anchor. (iii) Finally, in existing DAG protocols, an anchor requires at least 2 more DAG rounds to be committed; we refer to this as Anchor Commit Latency.
Key ideas.
@@
To reduce end-to-end latency, Shoal++ employs three techniques, respectively addressing each latency stage. First, Shoal++ identifies and optimizes an inefficiency in the existing Bullshark commit rule to reduce common case Anchor Commit Latency to only 4md. Next, we configure Shoal++ to dynamically attempt to make every node an anchor, thus eliminating Anchoring Latency. Shoal++ builds on Shoal’s idea to dynamically re-interpret the anchor assignment at each commit, and introduces small timeouts to help replicas advance rounds in lock-step to ensure that future rounds sufficiently reference anchor candidates. Finally, to minimize Queuing latencies, Shoal++ orchestrates multiple staggered DAG instantiations in parallel, and intertwines their resulting logs; this allows transactions that miss inclusion to simply board the new round of the next, staggered DAG.
@@
Put together, these techniques allow Shoal++ to reduce expected end-to-end latency to only 4.5md, shaving off 6md compared to Shoal, the state of the art DAG-BFT.
@@
We’ve implemented a prototype of Shoal++ and evaluated it empirically against two popular DAG-BFT protocols (Bullshark (Spiegelman et al., 2022a) and Shoal (Spiegelman et al., 2023)), a concurrent DAG-BFT protocol (Mysticeti (Babel et al., 2023)), and a traditional BFT protocols (Jolteon (Gelashvili et al., 2022)). Our results are promising, demonstrating matching throughput of existing DAG BFT solutions while reducing latency by up to 40% over Mysticeti and by up to 60% over Bullshark.
2. Model
@@
We adopt standard BFT assumptions (Castro and Liskov, 2002; Yin et al., 2019). We refer to participants that follow the protocol as correct, and those that deviate as faulty (or Byzantine). There are a total of $n=3f+1$ replicas, of which at most $f$ may be faulty at any time. We make no assumptions on the number of faulty clients. We assume the presence of a strong, yet static adversary that may arbitrarily delay messages and coordinate all faulty particpiants, but may not break standard cryptographic primitives such as signatures. We denote signed messages using $\langle m\rangle\sigma$. We assume that all signatures and Quorums are validated, and henceforth omit explicit mention.
@@
The goal of BFT consensus is to establish a common totally ordered log across all correct replicas. We operate under the partially synchronous model (Dwork et al., 1988). The system is safe at all times – no two correct replicas ever disagree on a committed prefix of the log. Liveness is not guaranteed during periods of asynchrony, but eventually there arrives a Global Stabilization Time (GST) after which the network (temporarily) behaves synchronously, and progress is guaranteed.
3. Dissecting DAG-BFT
@@
Before we present Shoal++, we provide some brief background on the core mechanisms of DAG-BFT, and analyze the latency breakdown of existing protocols. We focus implicitly on certified DAG protocols derived from Narwahl(Danezis et al., 2022). While some recent works propose also uncertified DAGs in an effort to save latency, this comes at a reduced robustness and introduces new sources of unwanted latency (§3.2).
3.1. DAG core
@@
At the core of DAG-BFT is the round-based, certified DAG structure proposed in Narwahl (Danezis et al., 2022). Replicas proceed through a series of structured rounds, in which replicas continuously propose new batches of transactions. Each replica maintains their own local view of a DAG.
@@
A replica adds another replica’s proposal as node in its local DAG once the proposal is certified. Proposals in round $r$ must reference $n-f$ certified proposals from round $r-1$ to be considered valid candidates, and each replica may certify at most one candidate per round. To certify a proposal for round a replica follows a simple Reliable Broadcast procedure.
  1. (1)
@@
  2. (2)
    Upon receiving a valid proposal, a replica checks whether it has already received a different proposal from the same author for round $r$. If it has, it ignores the proposal; if it has not, it casts a signed vote $S\coloneqq\langle r,d=hash(P)\rangle\sigma$ back to the proposing replica.
  3. (3)
@@
Figure 1 illustrates a DAG of certified nodes as it appears in a local view of one of the validators. We note that, because proposals are certified, no replicas can produce two conflicting nodes for the same (round, replica) position. All replicas thus eventually converge on the same local view of the DAG.
Figure 1. Narwhal’s round-based DAG from a validator’s local point of view. All nodes are certified.
3.1.1. Consensus core.
By itself, the DAG is no more than a scalable and robust mempool. It continues to grow at the pace of the network (i.e. responsively), and, by virtue of being certified, ensures that all disseminated transaction batches are reliably available. A second consensus layer is needed to establish agreement. It can either be modularly layered by running an external consensus blackbox that commits views of the DAG (e.g. Narwhal+Hotstuff (Danezis et al., 2022)), or logically embedded into the existing DAG messages without the need for additional messages (e.g. Tusk (Danezis et al., 2022), Bullshark (Spiegelman et al., 2022a), or Shoal (Spiegelman et al., 2023)). In the following we focus solely on embedded DAG-BFT protocols as they are more efficient.
Embedded consensus.
@@
Anchor nodes are deterministically chosen in advance (much like leaders in traditional BFT protocols) and placed at regular intervals. Bullshark (Spiegelman et al., 2022a), for instance, uses a round-robin scheme to select anchor candidates, and places one anchor candidate at every other round of the DAG. Figure 2 illustrates the anchor commit logic in Bullshark.
Figure 2. Bullshark Commit example. Replica 4 observes two ($f+1$) votes for anchor A1, and thus directly commits it. Replica 1 observes only one ($<f+1$) vote, and thus cannot directly commit A1. Upon committing A2, however, it observes that A1 is present in its causal history and indirectly commits A1, before ordering A2.
@@
To reduce the expected latency, Shoal (Spiegelman et al., 2023) optimizes Bullshark in two ways. First, it leverages the deterministic anchor commit sequence to dynamically re-interpret the anchor schedule upon every commit. Rather than placing anchors every other round, this allows it to place an anchor every round. Second, it introduces an anchor reputation scheme to attempt to select as candidates only the fastest, and best-connected replicas; such candidates are more likely to be included as edges. This allows Shoal to avoid timeouts in all but extremely unlikely scenarios.
3.2. Latency breakdown
While the design of embedded DAG-BFT protocols is appealingly simple, existing protocols pay a high latency premium. In the following, we breakdown different sources of latencies present in DAG-protocols, and analyze their respective costs in Bullshark and Shoal.
We denote as the e2e commit latency of a transaction the time between a replica first receives the transaction, and the time that the transaction is ordered. We separate the e2e commit latency of certified DAG-BFT protocols into three components:
  1. (1)
    Queuing latency: the time it takes for a transaction to be proposed.
  2. (2)
    Anchoring latency: the time it takes for a non-anchor node’s proposal to be ”picked up” by an eventually committing anchor.
  3. (3)
@@ 3.2. Latency breakdown
Anchoring latency.
A proposal becomes ”anchored” once it is referenced by an eventually committing anchor. Anchoring latency depends on two factors: (i) the frequency of anchor candidates, and (ii) the reliability with which said anchor candidates are committed (i.e. not skipped). [^2] The proposals of non-anchor candidates are subject to both factors; the proposals of anchors candidates themselves, trivially, are subject only to the latter. Bullshark has anchor candidates in every odd numbered round (only every second round). Consequently, the best-case inclusion latency of non-anchor proposals in even rounds is 1 round (3md), and 2 rounds (6md) in odd rounds; a best-case average of 4.5md. Shoal increases candidate frequency to every round, and thus improves the inclusion latency of non-anchor proposals to one round for all (3md). The practical reliability of commitment cannot be easily quantified by pen-and-paper, and is best left to empiric analysis as it depends on network configuration and message delivery patterns. Shoal demonstrates that the use of a reputation can aid significantly in ensuring that anchor candidates are reliably included (Spiegelman et al., 2023).
Anchor Commit latency.
@@
A note on un-certified DAGs.
A class of recent DAG-based designs (Keidar et al., 2022; Babel et al., 2023; Malkhi et al., 2023) proposes to replace nodes with un-certified proposals that are disseminated with best-effort. At first glance, such design holds promise by reducing latencies across the stack. However, removing certificates makes the DAG brittle, and can introduce a significant amount of unwanted synchronization on the critical path, and increase subsceptibility to timeout violations (Giridharan et al., 2024). Edges to uncertified, yet locally unavailable proposals must be validated by fetching missing data before advancing to avoid losing liveness in the face of Byzantine fabrications or unstable networks. In practice, such fetching can results in an additional 2md that may be necessary in each round; resulting, in sum, in latencies no better than in existing certified DAGs. In contrast, Shoal++ aims to demonstrates that certification is not the root-cause of high latencies, and that latency can be significantly reduced without sacrificing robustness.
4. Shoal++ Overview

### v2

# Shoal++: High Throughput DAG BFT Can Be Fast and Robust!

@@ Abstract
Today’s practical partially synchronous Byzantine Fault Tolerant consensus protocols trade off low latency and high throughput. On the one end, traditional BFT protocols such as PBFT and its derivatives optimize for latency. They require, in fault-free executions, only 3 message delays to commit, the optimum for BFT consensus. However, this class of protocols typically relies on a single leader, hampering throughput scalability. On the other end, a new class of so-called DAG-BFT protocols demonstrates how to achieve highly scalable throughput by separating data dissemination from consensus, and using every replica as proposer. Unfortunately, existing DAG-BFT protocols pay a steep latency premium, requiring on average 10.5 message delays to commit transactions.
@@
This work aims to soften this tension, and proposes Shoal++, a novel DAG-based BFT consensus system that offers the throughput of DAGs while reducing end-to-end consensus commit latency to an average of 4.5 message delays. Our empirical findings are encouraging, showing that Shoal++ achieves throughput comparable to state-of-the-art DAG BFT solutions while reducing latency by up to 60%, even under less favorable network and failure conditions.
**footnotetext: The work was done while the authors were interns at Aptos Labs.
1 Introduction
@@
This paper presents Shoal++, a novel partially synchronous DAG-based Byzantine Fault Tolerant (BFT) consensus protocol that matches the throughput of contemporary DAG-BFT protocols while significantly reducing latency, narrowing the gap with the theoretical optimum.
@@
BFT consensus offers the attractive abstraction of a single, always available server, that remains correct even as a subset of replicas may fail arbitrarily. Fueled by recent advances to simplicity and practicality [14, 46, 18, 34], interest and adoption of BFT protocols is skyrocketing: it is at the core of aspiring multi-national projects such as the digital Euro [2], confidential data sharing frameworks [5] and the growing Web3 industry. Already today, numerous blockchain companies [41, 39, 40, 43] are actively deploying BFT protocols [34, 33, 15, 30, 21], supporting millions of daily worldwide users.
@@
Following the seminal PBFT [14] protocol, most traditional BFT consensus protocols [46, 27, 23] operate in a primary-backup like regime in which a single, dedicated leader replica proposes an ordering of commands, and follower replicas vote to ensure consistency and durability.
@@
As demonstrated by PBFT, this approach can enjoy excellent latency in fault-free executions, requiring only three message delays (md) to order a leader’s proposal, the optimum achievable for BFT protocols. A dedicated leader, however, poses not only a single point of failure, but constrains the achievable throughput to the processing and networking bandwidth available to a single server. To improve scalability, and meet the high throughput demands of modern applications, recent proposals emphasize maximizing resource utilization and balancing load across all replicas.
Initial approaches propose multi-log frameworks [36] that partition the request space, and operate multiple consensus instances in parallel (each led by a different replica) whose logs are carefully intertwined. More recent designs take scalability a notch further by entirely separating data dissemination and consensus. Pioneered by Aleph [20] and DAG-Rider [25], and made practical by Narwahl [18] a new class of so-called DAG-BFT protocols have risen to popularity.
@@
State-of-the-art DAG protocols [34, 33] avoid running an explicit consensus protocol, and instead embed the consensus process onto the DAG directly: they ensure that all replicas converge on the same view of the DAG, and interpret it’s structure locally to implicitly determine commitment. Consensus simulates leaders by assigning pre-determined "fixpoint" nodes in the DAG (henceforth anchors) and confirms agreement of an anchor by waiting until it has gathered sufficient references in future DAG rounds that guarantee its durability.
@@
Upon committing an anchor node, all of its causally referenced proposals are implicitly committed and ordered. Such design is appealingly simple – it requires no additional communication for consensus and avoids the notoriously complex view change logic inherent to traditional BFT protocols. It is also very resource efficient – each replica acts as a proposer, thus allowing high throughput with a large number of replicas. As a result, DAG-BFT has prompted swift adoption by several large blockchain systems [43, 41].
@@
However, today’s DAG protocols fall short in terms of latency. They require several rounds, each consisting of reliable broadcasts to certify DAG nodes, to commit a transaction. Bullshark [34] for instance requires, even in fault-free cases, up to 4 rounds to commit a proposal (12md in expectation, §3.2); recent advances made by Shoal [33] reduce it to 3 rounds (10.5md expected, §3.2). A far cry from the 3md possible by traditional BFT protocols!
This work proposes Shoal++: a novel DAG-based BFT system that achieves near-optimal latency while maintaining the high throughput and robustness of the state-of-the-art certified DAG construction.
@@
We begin by analyzing sources of end-to-end (e2e) latency in DAG protocols (§3.2); at a high level, it can be broken down into three stages. (i) DAG proposals happen only at regular intervals (rounds), and transactions that narrowly miss inclusion (e.g. arrive just after a node is formed) must wait for the next round. We refer to the time a transaction must wait to be included in a DAG proposal as Queuing Latency. (ii) In order to be committed, a proposed DAG node must be referenced by a committed anchor. Unfortunately, in existing DAG protocols [34, 33], at most one anchor is scheduled per round [^1] Bullshark [34] schedules an anchor every second round; Shoal [33] improves it to every round. This results in additional Anchoring latency, the time until a node is referenced by the next committed anchor. (iii) Finally, in existing DAG protocols, an anchor requires at least two more DAG rounds to be committed; we refer to this as Anchor Commit Latency.
Key ideas
@@
To reduce end-to-end consensus latency, Shoal++ employs three techniques, respectively addressing each latency stage. First, Shoal++ identifies and optimizes an inefficiency in the existing Bullshark commit rule to reduce common case Anchor Commit Latency to only 4md. Next, we configure Shoal++ to dynamically attempt to make every node an anchor, thus eliminating Anchoring Latency. Shoal++ builds on Shoal’s idea to dynamically re-interpret the anchor assignment at each commit, and introduces small timeouts to help replicas advance rounds in lock-step to ensure that future rounds sufficiently reference anchor candidates. Finally, to minimize Queuing latencies, Shoal++ orchestrates multiple staggered DAG instantiations in parallel, and intertwines their resulting logs; this allows transactions that miss inclusion to simply board the new round of the next, staggered DAG.
@@
Put together, these techniques allow Shoal++ to reduce expected end-to-end consenus latency to only 4.5md, shaving off 6md compared to Shoal, the state of the art DAG-BFT.
@@
We’ve implemented a prototype of Shoal++ and evaluated it empirically against two popular DAG-BFT protocols (Bullshark [34] and Shoal [33]), a concurrent DAG-BFT protocol (Mysticeti [11]), and a traditional BFT protocols (Jolteon [21]). Our results are promising, demonstrating matching throughput of existing DAG BFT solutions while reducing latency by up to 60% over Bullshark in the common case and by up to 10x over Mysticeti in the failure case.
2 Model
@@
We adopt standard BFT assumptions [14, 46]. We refer to participants that follow the protocol as correct, and those that deviate as faulty (or Byzantine). There are a total of $n=3f+1$ replicas, of which at most $f$ may be faulty at any time. We make no assumptions on the number of faulty clients. We assume the presence of a strong, yet static adversary that may arbitrarily delay messages and coordinate all faulty particpiants, but may not break standard cryptographic primitives such as signatures. We denote signed messages using $\langle m\rangle_{\sigma}$. We assume that all signatures and Quorums are validated, and henceforth omit explicit mention.
@@
The goal of BFT consensus is to establish a common totally ordered log across all correct replicas. We operate under the partial synchrony model [19]. The system is safe at all times – no two correct replicas ever disagree on a committed prefix of the log. Liveness is not guaranteed during periods of asynchrony, but eventually there arrives a Global Stabilization Time (GST) after which the network (temporarily) behaves synchronously, and progress is guaranteed.
3 Dissecting DAG-BFT
@@
Before we present Shoal++, we provide a brief background on the core mechanisms of DAG-BFT, and analyze the latency breakdown of existing protocols. We implicitly focus on certified DAG protocols derived from Narwahl[18]. Although some recent works propose uncertified DAGs in an effort to save latency, this comes at the cost of reduced robustness and introduces new sources of unwanted latency (§3.2).
3.1 DAG core
@@
At the core of DAG-BFT is the round-based certified DAG structure proposed in Narwahl [18]. Replicas proceed through a series of structured rounds, in which replicas continuously propose new nodes containing batches of transactions. Each replica maintains their own local view of a DAG.
@@
A replica adds another replica’s proposal as node in its local DAG once the proposal is certified. Proposals in round $r$ must reference $n-f$ certified proposals from round $r-1$ to be considered valid candidates, and each replica may propose and certify at most one candidate per round. To certify a proposal for round a replica follows a simple Reliable Broadcast procedure.
  1. 1.
@@
  2. 2.
    Upon receiving a valid proposal, a replica checks whether it has already received a different proposal from the same author for round $r$. If it has, it ignores the proposal; if it has not, it casts a signed vote $S\coloneqq\langle r,d=hash(P)\rangle_{\sigma}$ back to the proposing replica.
  3. 3.
@@
Figure 1 illustrates a DAG of certified nodes as it appears in a local view of one of the validators. We note that because proposals are certified, no replicas can produce two conflicting nodes for the same (round, replica) position. Thus, all replicas eventually converge on the same local view of the DAG.
Figure 1: Narwhal’s round-based DAG from a validator’s local point of view.
3.1.1 Consensus core.
By itself, the DAG is no more than a scalable and robust mempool. It continues to grow at the pace of the network (i.e. responsively), and, by virtue of being certified, ensures that all disseminated transaction batches are reliably available. A consensus layer is needed to establish agreement. It can be modularly layered by running an external consensus blackbox that commits views of the DAG (e.g., Narwhal+Hotstuff [18]), or logically embedded into the existing DAG messages without the need for additional messages (e.g., Tusk [18], Bullshark [34], or Shoal [33]). The rest of the paper focuses on embedded DAG-BFT protocols as they are more efficient.
Embedded consensus
@@
Anchor nodes are chosen deterministically in advance (much like leaders in traditional BFT protocols) and placed at regular intervals. Bullshark [34], for instance, uses a round-robin scheme to select anchor candidates and places an anchor candidate per every other round of the DAG. Figure 2 illustrates the anchor commit logic in Bullshark.
Figure 2: Bullshark Commit example. Replica 4 observes $f+1=2$ votes for anchor A1, and directly commits it. Replica 1 observes only one vote ($<f+1$), and cannot directly commit A1; it commits A1 indirectly upon committing A2 and its causal history, ordering A1 before A2.
@@
To reduce the expected latency, Shoal [33] optimizes Bullshark in two ways. First, it leverages the deterministic anchor commit sequence to dynamically re-interpret the anchor schedule upon every commit. Rather than placing anchors every other round, this approach allows for an anchor to be placed every round. Second, it introduces an anchor reputation scheme to attempt to select as candidates only the fastest and well-connected replicas; such candidates are more likely to be included as edges. This allows Shoal to avoid timeouts in all but extremely unlikely scenarios.
3.2 Latency breakdown
In the following, we breakdown different sources of latencies present in DAG-protocols, and analyze their respective costs in Bullshark and Shoal. We denote as the e2e consensus latency of a transaction the time between a replica first receives the transaction, and the time that the transaction is ordered. We separate the e2e latency of certified DAG-BFT protocols into three components:
  1. 1.
    Queuing latency: the time it takes for a transaction to be proposed in a DAG node.
  2. 2.
    Anchoring latency: the time it takes for a non-anchor node’s proposal to be "picked up" by an eventually committing anchor.
  3. 3.
@@ 3.2 Latency breakdown
Anchoring latency
A proposal becomes "anchored" once it is referenced by an eventually committing anchor. Anchoring latency depends on two factors: (i) the frequency of anchor candidates, and (ii) the reliability with which said anchor candidates are committed (i.e. not skipped). [^2] The proposals of non-anchor candidates are subject to both factors; the proposals of anchors candidates themselves, trivially, are subject only to the latter. Bullshark has anchor candidates in every odd numbered round (i.e., only every second round). Consequently, the best-case inclusion latency of non-anchor proposals in even rounds is 1 round (3md), and 2 rounds (6md) in odd rounds; a best-case average of 4.5md. Shoal increases candidate frequency to every round, and thus improves the inclusion latency of non-anchor proposals to one round for all (3md). The practical reliability of commitment cannot be easily quantified by pen-and-paper, and is best left to empiric analysis as it depends on network configuration and message delivery patterns. Shoal demonstrates that the use of a reputation can aid significantly in ensuring that anchor candidates are reliably included [33].
Anchor Commit latency
@@
3.3 A note on uncertified DAGs
A class of recent DAG-based designs [26, 11, 28] proposes to replace DAG nodes with uncertified proposals that are disseminated with best-effort. At first glance, such designs holds promise by reducing latencies across the stack. However, removing certificates makes the DAG brittle: it can introduce a significant amount of unwanted synchronization on the critical path and increase susceptibility to timeout violations [22]. Edges to uncertified, yet locally unavailable proposals must be validated by fetching missing data (at least 2 additional md per round) before advancing to avoid losing liveness in the face of Byzantine fabrications or unstable networks. In practice, we find that even occasional message drops can significantly degrade overall latency (§8.3). In contrast, Shoal++ demonstrates that certification is not the root-cause of high latencies, and that latency can be significantly reduced without sacrificing robustness.
4 Shoal++ Overview